please help me solvethis problem: complete the following proof. Given: 2x+5=11 prove: x=3 Proof:

Answer by richard1234(7193)

Relying upon the axioms of arithmetic (e.g. a = b implies a + c = b + c, or a/c = b/c), we can find x:

2x + 5 - 5 = 11 - 5

2x = 6

x = 3